# Phase Matching Free Sensing with Undetected Light Using a Nonlinear Metasurface

### v1

@@ 0.0.2 Sensing through opaque media
We compare the theoretical prediction to the experimental visibility in Fig. 2(a) with a maximum of 55.0% $\pm$ 11.2%. Here, the visibility is calculated by fitting $N=A(1+Vcos(\phi))$ to the interference at each pixel on the camera, where $\phi$ is the phase and A is the DC intensity. As an interferometric technique, the corresponding phase map, Fig. 2(c), is simultaneously extracted. The saddle-like phase pattern is characteristic of interference between two fields; one diverging and the other converging. Ultimately, this condition resulting in maximised visibility is due to the same refractive lens focusing both the pump and seed to different focal points.
@@
Figure 2: Visibility and phase signal interference with and without a 240 µm Si window in the seed (idler) arm of the interferometer. Experimental visibility at the camera with no additional material in the seed (idler) path (a) and with the silicon window (b). The corresponding phase of interference, without (c), and with (d), the silicon window in the seed (idler) arm.
@@
An advantage of non-degenerate imaging and sensing with undetected light is the ability to observe at one wavelength while exploiting the optical properties of light-matter interaction at another. To highlight this effect, we use the same process of extracting visibility and phase information to interference data collected through an anti-reflection coated 240 µm piece of polished Si placed at the object plane. At 580 nm, the Si window is opaque, but at 1500 nm it is measured to have a transmission of 91.2 $\pm$ 0.9%, which leads to a theoretical maximum visibility of 52.3 $\pm$ 0.5% in Fig. S3(b) of the SI. In the experiment, Fig. 2(b), the visibility has dropped to a maximum of 24.4 $\pm$ 12.2%. The path length introduced by the Si window is $\sim 0.6$ mm; greater than the coherence length of interference. Re-establishing temporal overlap is likely to have reduced nonlinear emission on the second pass, and in turn, reduced visibility. A further consequence of this realignment is that a direct measure of the phase change cannot be taken from Fig. 2(c) and (d). However, the clear pattern in (d) shows phase information is retrieved despite the reduced visibility. Additionally, the change in phase pattern suggests the focal points of the two passes have reversed from the alignment in (c).
0.0.3 Effect of metasurface optical properties and single-mode operation.
@@
The derived visibility is dependent on the optical properties of the nonlinear metasurface. Visibility as a function of $\eta$, the ratio of nonlinear emission ($\sigma_{-}$/$\sigma_{+}$), shown in Fig. 3(a), peaks at $\eta$ = 0.67, offset from 1 due to a trade-off between the two sets of interfering fields in Eq. (5). The first interfering term depends only on Tf while the second depends on both Tf and Rf. Therefore, maximising Tf brings the upper limit of visibility down in Fig. 3(b), while bringing the optimal $\eta\mapsto 1$. Consequently, further optimisation could be achieved by iteratively tuning $\eta$ and Tf (Rf). The ratio of total FWM (signal) emission on the second pass compared to the first is represented by the balancing term $\kappa$. As both interfering terms in Eq. (5) depend on emission from each pass, visibility is maximised at $\kappa=1$ in Fig. 3(c). Fixing the optical parameters at the experimental and optimised values in Fig. 3(a-c) leads to the visibility curves in Fig. 3(d) as functions of the seed (idler) transmission. With this being said, an optimised metasurface may be difficult to engineer, requiring perfect transmission at the pump and seed wavelengths while balancing the bidirectionality of nonlinear emission with the reflection and transmission of FWM (signal).
@@
Figure 3: Visibility as functions of the average optical properties of the metasurface and transmission in the seed (idler) arm. Theoretical dependence of visibility on (a) $\eta$, (b) Tf, (c) $\kappa$ and (d) seed (idler) transmission, $T$. The experimentally measured and optimised values are marked by the vertical lines. Constant parameters are held at the experimentally measured values in Table I of the SI. (e) Experimental average visibility across three rows of pixels, at knife edge positions that reduce seed (idler) transmission at the object plane. (f) Experimental normalised average visibility across all pixels as a function of transmission at the object plane as the knife edge is introduced. The theoretical curve uses experimentally measured optical properties as constant parameters.
@@
In the stimulated FWM process at the metasurface the pump and seed modes generate a single idler mode. The resulting interferometer is therefore a single mode sensor and does not offer imaging capability. Any additional modes created by diffraction and/or refraction by an object do not interfere with the single mode of the first pass. By scanning a knife edge across the object plane of the interferometer we can measure a reduction of visibility, shown in Fig. 3(e). The profile of the FWM mode remains the same shape with horizontal visibilities falling together as the seed (idler) is partially to fully obstructed. Fig. 3(f) displays the normalised visibility averaged across the entire camera sensor area, for the same experiment. The knife edge position has been converted to normalised seed (idler) transmission, $T$, to identify the single mode object transmission. The theoretical model predicts the visibility using the experimentally measured optical parameters of the metasurface (Table I of the SI). The visibility is normalised to account for averaging across the camera sensor.
0.0.4 Spectral resolution and dispersion
@@
Figure 4: 2D FWM interference spectrograms for wavelength against pump-seed delay position. (a) Intensity of the FWM (signal) DC component indicating the chirp of spectral components. (b) Spectrally resolved FWM (signal) visibility as a function of pump-seed delay, measured by scanning the interferometer idler mirror at each delay position. (c) The spectrally resolved accumulative phase from all beams as a function of pump-seed delay showing the small intrinsic dispersion of the interferometer. For (b) and (c), data extracted with a coefficient of determination, R2, < 0.7 has been removed.
@@
Despite the clear peak in FWM (signal) emission in Fig. 4(a) the visibility in Fig. 4(b) shows a flatter profile, and a continuous phase relationship can be extracted across the entire FWM (signal) spectral range in Fig. 4(c).
@@
The spectral shift in Fig. 4(c) shows the difference in phase between FWM (signal) generated at the first and second pass. Therefore, the phase shift comes from the combined dispersion experienced by the seed (idler), pump, and FWM (signal) after the first pass. All three pulses travel twice through the 175 $\pm$ 13 µm borosilicate substrate of the metasurface, while the seed (idler) also travels twice through a 1 mm thick dichroic mirror nominally positioned at 45 degrees to the incoming beams. Averaging the phase shift in Figure 4 (c) across all delay positions for 575 to 585 nm gives $-1.56\pm 0.29$. Importantly, the interferometer is only sensitive to variations in the object’s phase, while changes in the FWM phase for different delay positions have no effect on the phase variation, as seen in Fig. 4(c). Interestingly, a Fourier transform of the interferometer visibility and phase would reveal ultrafast transients of the object transmission with a resolution less than the pulse duration.
@@
In conclusion, we have demonstrated the operation of a nonlinear interferometer that uses a nonlinear plasmonic metasurface as a source to generate signal and idler photons through stimulated FWM to implement classical sensing with undetected light. The ability to pass amplitude and phase information from the NIR (1500 nm) to the visible (580 nm) is demonstrated by sensing through a 240 µm Si window while detecting the emission and its phase on a Si camera. Meanwhile, a knife edge was used to modulate the single-mode transmission of the interferometer. A theory of the interferometer based on any thin nonlinear source with intrinsic reflections accurately predicts the experimental visibility as a function of NIR seed (idler) transmission.
@@
The single mode nature of interference—brought about by seeding the FWM (signal) with single mode lasers—provides no spatial information for imaging. However, additional modes could be generated before the first pass through the metasurface. Alternatively, imaging could be achieved by adapting the interferometer to operate in real space, where the object plane is imaged onto the metasurface, which is in turn imaged onto the camera.
@@
The spectral breadth of the signal and idler generated through pulsed illumination allows spectrally resolved visibility and phase to be measured, assessing dispersion within the interferometer. Additionally, pulsed illumination opens up avenues for ultra-fast pump-probe spectroscopy with undetected photons, such as lifetime measurements for molecular excited states and vibrations.
@@
In combination, pulsed illumination and stimulated nonlinear emission reduce the requirement for millimetre scale exotic crystals and phase matching. The nonlinear interferometer system introduced here has the potential to reach any wavelength range where a laser and metasurface resonances are spectrally overlapped. Therefore, operation deep in the MIR ($\sim$10 µm) is conceivable, where micron scale resonators are easy to fabricate and pulsed lasers are readily available.
@@ Methods
Optical measurements were taken using a seed laser at 1500 nm ($\leq$ 228 fs) generated by a Coherent Chameleon Compact OPO, pumped by a Coherent Chameleon Ultra II at 835 nm ($\leq$ 146 fs). The seed was sent to a piezo-controlled delay line before being combined with the pump at a 1 µm short-pass dichroic mirror (Thorlabs DMSP1000). Both beams were passed through a ×$\times 4$ reflective beam expander (Thorlabs BE04R) and focused onto the sample using a 20 cm focal length refractive lens. After the sample, the transmitted pump and forward generated FWM (signal) were separated from the seed (signal) by a 1 µm long-pass dichroic mirror (Thorlabs DMLP1000). Identical curved mirrors, each with a 20 cm radius of curvature, reflected the respective beams back to the sample. The pump and FWM (signal) path lengths were varied by a piezo-controlled stage connected to the curved mirror in their path. A 650 nm long-pass dichroic mirror (Thorlabs DMLP650) isolated the FWM (signal) and directed it onto an EM-CCD camera (Hamamatsu C9100-23B ImagEM X2).
For spectrally resolved measurements, the camera was replaced with a 20 cm focal length lens to collimate the FWM (signal). A pinhole was used to select the centre of the interference pattern that was subsequently focused onto the entrance slit of the spectrometer (Acton SP2300) and detected on a Si-based CCD.
@@
Contains a detailed derivation of theory, relevant optical parameters of the metasurface, and predicted results based on these. A video of the interference pattern as the phase is shifted is also included.
@@ Acknowledgments
T.S.M. acknowledges the support of the UK government Department for Science, Innovation and Technology through the UK National Quantum Technologies Programme and the National Physical Laboratory. This work is supported by the UK Quantum Technology Hubs in Sensing Imaging and Timing EP/Z533166/1 and Quantum Imaging EP/T00097X/1 and by the UK Engineering and Physical Sciences Research Council (EPSRC), through the Reactive Plasmonics EP/M013812/1 and Catalysis Plasmonics EP/W017075/1 Programme Grants. A.S.C acknowledges support from The Royal Society (URF/R/221019, RF/ERE/210098, RF/ERE/221060).
@@
M.R. fabricated the metasurface. T.S.M. took optical measurements. R.F.O., T.S.M. and N.G. developed the theory. T.S.M. wrote the paper with contributions from all co-authors. R.F.O., N.G., A.S.C., and C.C.P. conceived and supervised the project.

### v2

@@ 0.0.2 Sensing through opaque media
We compare the theoretical prediction to the experimental visibility in Fig. 2(a) with a maximum of 55.0% $\pm$ 1.7%. Here, the visibility is calculated by fitting $N=D(1+V\mathrm{cos}(\phi))$ to the interference at each pixel on the camera, where $\phi$ is the phase and $D$ is the DC intensity. As an interferometric technique, the corresponding phase map, Fig. 2(c), is simultaneously extracted. The saddle-like phase pattern is characteristic of interference between two fields; one diverging and the other converging. Ultimately, this condition resulting in maximised visibility is due to the same refractive lens focussing both the pump and seed to different focal points.
@@
Figure 2: Visibility and phase of the FWM (signal) interference, centred at 580 nm, with and without a 240 µm silicon window in the seed (idler) arm of the interferometer. Experimental visibility at the camera with no additional material in the seed (idler) path (a) and with the silicon window (b). The corresponding phase of interference, without (c), and with (d), the silicon window in the seed (idler) arm.
@@
An advantage of non-degenerate imaging and sensing with undetected light is the ability to observe at one wavelength while exploiting the optical properties of light-matter interaction at another. To highlight this effect, we apply the same process of extracting visibility and phase information to interference data collected through an anti-reflection coated 240 µm piece of polished silicon placed at the object plane. At 580 nm, the silicon window is opaque, but at 1500 nm it is measured to have a single pass transmission of 91.3 $\pm$ 0.9%, which leads to a theoretical maximum visibility of 52.3 $\pm$ 0.5% in Fig. S5(b) of the SI. In the experiment, Fig. 2(b), the visibility has dropped to a maximum of 24.4 $\pm$ 1.7%. The path length introduced by the silicon window is $\sim 1.2$ mm; greater than the coherence length of interference. Re-establishing temporal overlap is likely to have reduced nonlinear emission on the second pass, and in turn, reduced visibility. A further consequence of this realignment is that a direct measure of the phase change cannot be taken from Fig. 2(c) and (d). However, the clear pattern in (d) shows phase information is retrieved despite the reduced visibility. Additionally, the change in phase pattern suggests the focal points of the two passes have reversed from the alignment in (c).
0.0.3 The metasurface’s optical properties & single-mode operation.
@@
The derived visibility is dependent on the optical properties of the nonlinear metasurface. Visibility as a function of $\eta$, the ratio of nonlinear emission ($\sigma_{-}$$\sigma_{+}$), shown in Fig. 3(a), peaks at $\eta$ = 0.67, offset from 1 due to a trade-off between the two sets of interfering fields in Eq. (5). The first interfering term depends only on Tf while the second depends on both Tf and Rf. Therefore, maximising Tf brings the upper limit of visibility down in Fig. 3(b), while bringing the optimal $\eta\mapsto 1$. Consequently, further optimisation could be achieved by iteratively tuning $\eta$ and Tf (Rf). The ratio of total FWM (signal) emission on the second pass compared to the first is represented by the balancing term $\kappa$. As both interfering terms in Eq. (5) depend on emission from each pass, visibility is maximised at $\kappa=1$ in Fig. 3(c). Fixing the optical parameters at the experimental and optimised values in Fig. 3(a-c) leads to the visibility curves in Fig. 3(d) as functions of the seed (idler) transmission. With this being said, an optimised metasurface may be difficult to engineer, requiring perfect transmission at the pump and seed wavelengths while balancing the bidirectionality of nonlinear emission with the reflection and transmission of FWM (signal).
@@
Figure 3: Visibility as functions of the optical properties of the metasurface and transmission in the seed (idler) arm. Theoretical dependence of visibility on (a) $\eta$, (b) Tf and (c) $\kappa$. The experimentally measured and theoretically optimised metasurface parameters are marked by the solid and dashed vertical lines, respectively. Constant parameters are held at the values in Table SI of the SI. (d) Dependence of visibility on seed (idler) transmission, $T$, for the experimental measured (solid curve) and theoretically optimised (dashed curve) metasurface parameters. (e) Experimental average visibility measured at knife edge positions that reduce seed (idler) transmission. The knife edge is brought along the horizontal axis of the object plane and visibility is averaged across three rows of pixels. (f) Experimental visibility as the knife edge is introduced, averaged across all pixels and normalised to the unobstructed visibility. The theoretical curve uses the experimentally measured optical properties.
@@
In the stimulated FWM process at the metasurface, the pump and seed modes generate a single idler mode. The resulting interferometer is therefore a single-mode sensor and does not offer imaging capability. Any additional modes created at an object do not interfere with the single mode of the first pass. By scanning a knife edge horizontally across the object plane of the interferometer we can measure a reduction in visibility across the horizontal (x-axis) of the camera, as shown in Fig. 3(e). The knife edge starts by not obstructing the seed (idler) beam at 0.00 mm to fully obstructing it at 6.25 mm, and visibility is calculated at each knife edge position. The profile of the FWM mode remains the same shape, with horizontal visibilities falling together as the seed (idler) is partially to fully obstructed. Fig. 3(f) displays the normalised visibility averaged across the entire camera sensor area for the same experiment. The knife edge position has been converted to normalised seed (idler) transmission, $T$, to identify the single-mode object transmission. The theoretical model predicts visibility using the experimentally measured optical parameters of the metasurface (Table I of the SI). The visibility is normalised to account for averaging across the camera sensor.
0.0.4 Spectroscopy & dispersion
@@
[…]
@@
Despite the clear peak in FWM (signal) emission in Fig. 4(a), the visibility in Fig. 4(b) shows a flatter profile, and a continuous phase relationship can be extracted across the entire spectral mode.
@@
The spectral shift in Fig. 4(c) shows the difference in phase between FWM (signal) generated at the first and second pass. Therefore, the phase shift comes from the combined dispersion experienced by the seed (idler), pump, and FWM (signal) after the first pass. All three pulses travel twice through the 175 µm borosilicate substrate of the metasurface, while the seed (idler) also travels twice through a 1 mm thick dichroic mirror nominally positioned at 45 degrees to the incoming beams. Averaging the phase shift in Fig. 4(c) across all delay positions from 575 to 585 nm gives $1.56\pm 0.29$. Importantly, the interferometer is only sensitive to variations in the object’s phase, while changes in the FWM (signal) phase for different delay positions have no effect on the phase variation, as seen in Fig. 4(c). Interestingly, with the addition of a second pulsed beam to illuminate the object plane, a Fourier transform of the interferometer visibility and phase would reveal ultrafast transients of the object transmission with a resolution less than the pulse duration.Kaur2024PulsedQFTIR With this in mind, the pulsed modes are expected to preserve their transform limited nature in our system; experiencing no temporal broadening from group velocity mismatch or phase matching bandwidth narrowing associated with bulk nonlinear material.
@@
In conclusion, we have demonstrated the operation of a nonlinear interferometer that uses a nonlinear plasmonic metasurface as a source to generate signal and idler photons through stimulated FWM to implement classical sensing with undetected light. The ability to pass amplitude and phase information from the NIR (1500 nm) to the visible (580 nm) is demonstrated by sensing through a 240 µm silicon window while detecting the emission and its phase on a silicon camera. Meanwhile, a knife edge was used to modulate the single-mode transmission of the interferometer. A theory of the interferometer based on any thin nonlinear source with intrinsic reflections accurately predicts the experimental visibility as a function of NIR seed (idler) transmission.
@@
The single-mode nature of interference – brought about by seeding the FWM (signal) with single mode lasers – provides no spatial information for imaging. However, additional modes could be generated before the first pass through the metasurface. Alternatively, imaging could be achieved by adapting the interferometer to operate in real space, where the object plane is imaged onto the metasurface, which is in turn imaged onto the camera. Similarly, for spectroscopy, the system can utilise the broad spectral range of the resonant metasurface, either by tuning the seeding wavelength or illuminating with a broadband seed beam.
@@
The spectral breadth of the signal and idler generated through pulsed illumination allows for spectrally resolved visibility and phase to be measured, assessing dispersion within the interferometer. Additionally, pulsed illumination opens up avenues for ultra-fast pump-probe spectroscopy with undetected photons, such as lifetime measurements for molecular excited states and vibrations.
@@
In combination, pulsed illumination and stimulated nonlinear emission reduce the requirement for millimetre-scale exotic crystals and phase matching. The nonlinear interferometer system introduced here has the potential to reach any wavelength range where a laser and metasurface resonances are spectrally overlapped. Therefore, operation deep in the MIR ($\sim$10 µm) is conceivable, where micron-scale resonators are easy to fabricate and pulsed lasers are readily available. Additionally, the interferometer model developed is independent of the thin-film source used, including thin films of nonresonant nonlinear crystals, allowing the operating wavelengths to be defined solely by the pump and seed beams.
@@ Methods
Optical measurements were taken using a seed laser at 1500 nm ($\lesssim$ 228 fs) generated by a Coherent Chameleon Compact OPO, pumped by a Coherent Chameleon Ultra II at 835 nm ($\lesssim$ 146 fs). The seed was sent to a piezo-controlled delay line before being combined with the pump at a 1 µm short-pass dichroic mirror (Thorlabs DMSP1000). Both beams were passed through a $\times 4$ reflective beam expander (Thorlabs BE04R) and focused onto the sample using a 20 cm focal length refractive lens. After the sample, the transmitted pump and forward generated FWM (signal) were separated from the seed (signal) by a 1 µm long-pass dichroic mirror (Thorlabs DMLP1000). Identical curved mirrors, each with a 20 cm radius of curvature, reflected the respective beams back to the sample. The pump and FWM (signal) path lengths were varied by a piezo-controlled stage connected to the curved mirror in their path. A 650 nm long-pass dichroic mirror (Thorlabs DMLP650) isolated the FWM (signal) and directed it onto an EM-CCD camera (Hamamatsu C9100-23B ImagEM X2). FWM (signal) emission and interference were recorded using $800$ ms exposure time and an EM gain of $600\times$.
For spectrally resolved measurements, the camera was replaced with a 20 cm focal length lens to collimate the FWM (signal). A pinhole was used to select the centre of the interference pattern that was subsequently focused onto the entrance slit of the spectrometer (Acton SP2300) and detected on a silicon-based CCD.
@@
The supporting information contains; experimentally measured optical parameters of the metasurface, a more detailed derivation of the theory included in this article, the theory behind two additional interferometer configurations, and calculations of predicted visibility for our system. A video of the interference pattern as the phase is varied is also included.
@@ Acknowledgments
T.S.M. acknowledges the support of the UK government Department for Science, Innovation and Technology through the UK National Quantum Technologies Programme and the National Physical Laboratory. This work is supported by the UK Quantum Technology Hubs in Sensing Imaging and Timing EP/Z533166/1 and Quantum Imaging EP/T00097X/1 and by the UK Engineering and Physical Sciences Research Council (EPSRC), through the Reactive Plasmonics EP/M013812/1 and Catalysis Plasmonics EP/W017075/1 Programme Grants. M.R. acknowledges support from the UK Research and Innovation Future Leaders Fellowship (MR/T040513/1). A.S.C. acknowledges support from The Royal Society (URF/R/221019, RF/ERE/210098, RF/ERE/221060).
@@
M.R. fabricated the metasurface. T.S.M. collected optical measurements. R.F.O., T.S.M. and N.G. developed the theory. T.S.M. wrote the paper with contributions from all co-authors. R.F.O., N.G., A.S.C. and C.C.P. conceived and supervised the project.